# Holmes-VAU: Towards Long-term Video Anomaly Understanding at Any Granularity

### v1

@@ 2 Related Works
Hierarchical Video Understanding. Video understanding is a challenging task due to its temporal-scale diversity. To better comprehend videos, many previous works have focused on both datasets and models in hierarchical video understanding. For example,  [24, 43, 32, 3] proposed fine-grained action recognition and localization datasets,  [16] provided free-form hierarchical captions for hour-long videos at multiple temporal scales, and  [18, 41, 8, 57, 66] trained models on hierarchical levels to obtain better video feature representation. Recently, to assess the capability of video vision-language models (VLMs) in handling challenges in real-world scenarios, several video benchmarks [12, 10] have been built, incorporating data at multiple temporal scales for evaluation. Unlike these works, we dive into the field of Video Anomaly Understanding, thus filling the gap in multi-scale annotations in this area.
@@ 3.2 LLM-Empowered Data Engine
Hierarchical Free-text Annotation. To fully extract semantic information from the clip-level videos, we utilize a powerful off-the-shell video perception model LLaVA-Next-Video [28] to generate detailed captions for each clip. We also include the UCA dataset  [61], which provides manually annotated captions for video clips in the UCF-Crime  [44] dataset. Then, we use an LLM [2] to consolidate all clip captions within an event, generating an event-level video summary. Specifically, we design prompts [^1] For detailed prompts, please refer to the appendix. to guide the LLM to produce three parts of content for each anomalous event summary: 1) Judgment: A determination of whether an anomaly exists and its specific category, 2) Description: A detailed description of the anomalous or normal event, 3) Analysis: The reasoning behind the anomaly judgment, including causal analysis. To guide the LLM in generating reliable responses, we also inject the event’s category label (e.g., ”Shooting”, ”Explosion”) into the LLM prompt. We consolidate all event-level summaries to obtain the video-level summary. This results in free-text annotations across multiple temporal scales, including short-term visual perception (clip-level) and long-term anomaly reasoning (event-level, video-level).
@@ 5.1 Experiment Setup
Metric. We assess the anomaly understanding ability from two aspects: anomaly detection and reasoning. 1) For anomaly detection, we use the anomaly scores output by the Anomaly Scorer as the prediction and perform the evaluation. Following  [14, 44, 71, 64], we use AUC and AP to quantify detection performance, which is evaluated only on the video level. 2) For anomaly reasoning, we annotate instruction data from the UCF-Crime and XD-Violence test sets, which have been carefully reviewed and filtered by annotators. We finally collected 3,300 test samples at multiple granularities. We calculate metrics including BLEU [39], CIDEr [49], METEOR [4] and ROUGE [26] to measure the quality of the reasoning text output by the model, comparing with the annotated ground truth text.
@@
Table 1: Comparison of detection performance with state-of-the-art Video Anomaly Detection approaches. We include the results of explainable and non-explainable methods. $\ast$” represents the result reported in  [64].
@@
Conv-AE [14] - 27.25 50.60
GODS [50] I3D N/A 70.46
GCL [63] ResNext N/A 71.04
DYANNET [46] I3D N/A 84.50
MIST [11] I3D N/A 82.30
Wu et al. [53] I3D 78.64 82.44
RTFM [47] I3D 77.81 84.30
MSL [22] I3D 78.28 85.30
S3R [52] I3D 80.26 85.99
MGFN [5] I3D 79.19 86.98
UR-DMU [71] I3D 81.66 86.97
CLIP-TSA [17] ViT 82.19 87.58
VadCLIP [55] ViT 84.51 88.02
Yang et al. [59] ViT 83.68 87.79
Wu et al. [54] ViT 66.53 86.40
@@ 5.4 Qualitative Comparision
We provide qualitative comparisons between Holmes-VAU and existing MLLMs in Fig 4. The results demonstrate that Holmes-VAU can accurately identify anomalies in videos and provide accurate and complete explanations, highlight the effectiveness and advantage of Holmes-VAU in perceiving video events and analyzing anomalies
@@ 6 Conclusion
Limitations and Future Work. While our work demonstrates significant strides in multi-granular video anomaly understanding, several limitations present opportunities for future enhancement. First, optimizing for real-time streaming remains a challenge; our sparse sampling approach improves efficiency, but further refinement is necessary for seamless deployment in streaming contexts. Additionally, our work has so far focused on surveillance data; extending our framework to other domains, such as industrial monitoring and medical diagnostics, will help validate its generalization capabilities. Lastly, integrating additional sensory data, like audio, and scalable hierarchical annotation could enhance anomaly detection and broaden applicability.
Acknowledgement This work is supported by the National Natural Science Foundation of China under grant U22B2053 and 623B2039.
@@ A.2 Hierarchical Free-text Annotation
[linenos, frame=lines, breaklines]json ”video”: ”v=2rfyeR-YaJw__#1_label_G-0-0”, ”n_frames”: 1940, ”fps”: 24.0, ”label”: [”Explosion”], ”clips”: [[[5.583, 11.903], [11.903, 18.222], [18.222, 24.542]], [[36.167, 43.48], [43.48, 50.792]]], ”clip_captions”:[ [ ”A military tank moving across a barren landscape with low-rise buildings and sparse vegetation. the sky is overcast, and the overall color palette is muted with earthy tones.”, ”A series of images depicting a barren landscape with a few buildings in the background. the foreground consists of a rocky terrain with sparse vegetation. the sky is overcast, and there are no visible people or moving objects.”, ”A silhouette of a person operating a large, mounted weapon on a rocky terrain under a clear sky. the individual appears to be adjusting or aiming the weapon.” ], [ ”A series of images depicting a hazy, dusty environment with buildings in the background. the focus is on a red light source that appears to be moving or flashing, possibly indicating some form of activity or event occurring in the area.”, ”A series of images depicting a military tank moving across a landscape with buildings in the background. the tank is seen from different angles, and there are visible explosions occurring in the distance.” ] ], ”events”: [[5.583, 24.542], [36.167, 50.792]], ”event_summary”: [ ”The anomaly exists, specifically an Explosion anomaly. The basis for judging the anomaly is the presence of an unexpected and sudden release of energy, likely accompanied by a loud noise and visible effects such as smoke, fire, or debris.”, ”The anomaly exists and its specific name is Explosion. The anomaly event is a series of visible explosions occurring in the distance as a military tank moves across a landscape with buildings in the background. The basis for judging the anomaly is the sudden and intense release of energy and light in the form of explosions, which is an unusual and notable event in the otherwise desolate and quiet war-torn environment depicted in the video”], ”video_summary”: ”The anomaly exists, specifically named as Explosion. The anomaly event depicts a series of visible explosions occurring in the distance as a military tank moves across a landscape with buildings in the background, characterized by sudden and intense releases of energy and light, accompanied by loud noise and visible effects such as smoke, fire, or debris. The basis for judging the anomaly lies in the unusual and notable nature of these explosions, which stand out against the otherwise desolate and quiet war-torn environment depicted in the video, making them an unexpected and sudden release of energy that grabs attention.”
@@ A.3 Hierarchical Instruction Data Construction
    ”Describe the video briefly.”
@@
    ”Describe the main events that take place in this video.”
@@
    ”Give a short description of the video.”
@@
    ”What happened in this video?”
@@
    ”Generate a brief caption for the video.”
@@
    ”Can you provide a brief description of the video?”
@@
    ”Briefly describe the main subjects and their actions in the video.”
@@
    ”Provide a short overview of what happens in the video?”
@@
    ”Describe the key moments that showcase the subjects’ activities in the video.”
@@
    ”Describe the sequence of events involving the main subjects in the video.”
@@
    ”What activities happen throughout the video?”
@@
    ”Describe the main subjects and their roles in the video.”
@@
    ”What key moments stand out in the video?”
@@
    ”What are the primary activities showcased in the video?”
@@
    ”What happens to the main subjects as the video progresses?”
@@
    ”What is a brief overview of what happens in the video?”
@@
    ”Describe the main subjects and their contributions to the video.”
@@
    ”Describe the key events in the video.”
@@
    ”Describe the video’s main activities.”
@@
    ”Can you describe the main action in this video briefly?”
@@
    ”Describe the video clip concisely.”
@@
    ”Provide a brief description of the given video clip.”
@@
    ”Summarize the visual content of the video clip.”
@@
    ”Give a short and clear explanation of the subsequent video clip.”
@@
    ”What types of anomalies are shown in the video clip?”
@@
    ”Are there any anomaly events detected in the video?”
@@
    ”Detect and classify the anomaly events in the video.”
@@
    ”Identify any abnormal behaviors depicted in the video.”
@@
    ”Determine whether there are anomaly events in the video and the specific name of the anomaly.”
@@
    ”What anomalies can be identified in the video?”
@@
    ”What categories of anomalies can be found in the video?”
@@
    ”Could you point out any abnormal actions in the video?”
@@
    ”Point out the abnormal actions in the video.”
@@
    ”Are there anomalies observed in the video clip?”
@@
    ”Describe the anomaly events observed in the video.”
@@
    ”Could you describe the anomaly events observed in the video?”
@@
    ”Could you specify the anomaly events present in the video?”
@@
    ”Give a description of the detected anomaly events in this video.”
@@
    ”Could you give a description of the anomaly events in the video?”
@@
    ”Provide a summary of the anomaly events in the video.”
@@
    ”Could you provide a summary of the anomaly events in this video?””
@@
    ”What details can you provide about the anomaly in the video?”
@@
    ”How would you detail the anomaly events found in the video?”
@@
    ”How would you describe the particular anomaly events in the video?”
@@
    ”Why do you judge this event to be anomalous?”
@@
    ”Can you provide the reasons for considering it anomalous?”
@@
    ”Can you give the basis for your judgment of this event as an anomaly?”
@@
    ”What led you to classify this event as an anomaly?”
@@
    ”Could you provide the reasons for considering this event as abnormal?”
@@
    ”What evidence do you have to support your judgment of this event as an anomaly?”
@@
    ”Can you analyze the factors contributing to this anomalous event?”
@@
    ”Could you share your analysis of the anomalous event?”
@@
    ”What patterns did you observe that contributed to your conclusion about this event being an anomaly?”
@@
    ”How do the characteristics of this event support its classification as an anomaly?”
@@
[linenos, frame=lines, breaklines]json ”id”: ”Abuse007_x264_E0C2”, ”type”: ”clip”, ”video”: ”ucf-crime/clips/train/Abuse007_x264_E0C2.mp4”, ”conversations”: [ ”from”: ”human”, ”value”: ”Describe the key events in the video.”, ”from”: ”gpt”, ”value”: ”The man in shorts was slapped to the ground by the strong man opposite.” ]
@@
[linenos, frame=lines, breaklines]json ”id”: ”Abuse018_x264_E0”, ”type”: ”event”, ”video”: ”ucf-crime/events/train/Abuse018_x264_E0.mp4”, ”conversations”: [ ”from”: ”human”, ”value”: ”Identify any abnormal behaviors depicted in the video.”, ”from”: ”gpt”, ”value”: ”An anomaly exists, specifically an instance of Police Brutality.”, ”from”: ”human”, ”value”: ”Could you provide a summary of the anomaly events in this video?”, ”from”: ”gpt”, ”value”: ”The anomaly event involves the police officer kicking the man who rushed towards him, causing the man to fall heavily to the ground with his head on his back.”, ”from”: ”human”, ”value”: ”Could you share your analysis of the anomalous event?”, ”from”: ”gpt”, ”value”: ”The excessive use of force by the police officer, which deviates from standard protocol and may be considered an abuse of authority.” ]
@@
[linenos, frame=lines, breaklines]json ”id”: ”v=xLFbqPzSI2Y__#1_label_B4-0-0”, ”type”: ”video”, ”video”: ”xd-violence/videos/train/ v=xLFbqPzSI2Y__#1_label_B4-0-0.mp4”, ”conversations”: [ ”from”: ”human”, ”value”: ”Are there any anomaly events detected in the video?”, ”from”: ”gpt”, ”value”: ”The anomaly exists and its specific name is the Riot anomaly.”, ”from”: ”human”, ”value”: ”Could you give a description of the anomaly events in the video?”, ”from”: ”gpt”, ”value”: ”The anomaly event depicts a chaotic and intense urban unrest scene, starting with a chaotic street scene where individuals are running away from an area of fire and smoke, with people appearing to be in a state of panic or urgency, and some being sprayed with water by individuals in uniform. The scene escalates to show a growing fire in the background, police officers in riot gear standing on the street, and a formation of police officers spread out across the road, suggesting an ongoing riot or civil unrest. ”, ”from”: ”human”, ”value”: ”Can you give the basis for your judgment?”, ”from”: ”gpt”, ”value”: ”The presence of multiple indicators of a riot or emergency situation, including fire, smoke, and people fleeing in panic, as well as the sudden appearance of flames, smoke, and police in riot gear, and the presence of violent and destructive events, such as fires and explosions, and the response of authorities, including the use of tear gas or pepper spray, which deviate significantly from normal, peaceful urban activity.” ]

### v2

@@ 2 Related Works
Hierarchical Video Understanding. Video understanding is a challenging task due to its temporal-scale diversity. To better comprehend videos, many previous works have focused on both datasets and models in hierarchical video understanding. For example,  [24, 45, 32, 3, 54, 35] proposed fine-grained action recognition and localization datasets,  [16] provided free-form hierarchical captions for hour-long videos at multiple temporal scales, and  [18, 43, 8, 60, 69, 38] trained models on hierarchical levels to obtain better video feature representation. Recently, to assess the capability of video vision-language models (VLMs) in handling challenges in real-world scenarios, several video benchmarks [12, 10] have been built, incorporating data at multiple temporal scales for evaluation. Unlike these works, we dive into the field of Video Anomaly Understanding, thus filling the gap in multi-scale annotations in this area.
@@ 3.2 LLM-Empowered Data Engine
Hierarchical Free-text Annotation. To fully extract semantic information from the clip-level videos, we utilize a powerful off-the-shell video perception model LLaVA-Next-Video [28] to generate detailed captions for each clip. We also include the UCA dataset  [64], which provides manually annotated captions for video clips in the UCF-Crime  [46] dataset. Then, we use an LLM [2] to consolidate all clip captions within an event, generating an event-level video summary. Specifically, we design prompts [^1] For detailed prompts, please refer to the appendix. to guide the LLM to produce three parts of content for each anomalous event summary: 1) Judgment: A determination of whether an anomaly exists and its specific category, 2) Description: A detailed description of the anomalous or normal event, 3) Analysis: The reasoning behind the anomaly judgment, including causal analysis. To guide the LLM in generating reliable responses, we also inject the event’s category label (e.g., "Shooting", "Explosion") into the LLM prompt. We consolidate all event-level summaries to obtain the video-level summary. This results in free-text annotations across multiple temporal scales, including short-term visual perception (clip-level) and long-term anomaly reasoning (event-level, video-level).
@@ 5.1 Experiment Setup
Metric. We assess the anomaly understanding ability from two aspects: anomaly detection and reasoning. 1) For anomaly detection, we use the anomaly scores output by the Anomaly Scorer as the prediction and perform the evaluation. Following  [14, 46, 74, 67], we use AUC and AP to quantify detection performance, which is evaluated only on the video level. 2) For anomaly reasoning, we annotate instruction data from the UCF-Crime and XD-Violence test sets, which have been carefully reviewed and filtered by annotators. We finally collected 3,300 test samples at multiple granularities. The test set contains 2200/732/398 samples at clip/event/video levels. We calculate metrics including BLEU [41], CIDEr [51], METEOR [4] and ROUGE [26] to measure the quality of the reasoning text output by the model, comparing with the annotated ground truth text.
@@
Table 1: Comparison of detection performance with state-of-the-art Video Anomaly Detection approaches. We include the results of explainable and non-explainable methods. $\ast$" represents the result reported in  [67].
@@
Conv-AE [14] (CVPR’16) - 27.25 50.60
GODS [52] (ICCV’19) I3D N/A 70.46
GCL [66] (CVPR’22) ResNext N/A 71.04
DYANNET [48] (WACV’23) I3D N/A 84.50
MIST [11] (CVPR’21) I3D N/A 82.30
Wu et al. [56] (ECCV’20) I3D 78.64 82.44
RTFM [49] (ICCV’21) I3D 77.81 84.30
MSL [22] (AAAI’22) I3D 78.28 85.30
S3R [55] (ECCV’22) I3D 80.26 85.99
MGFN [5] (AAAI’23) I3D 79.19 86.98
UR-DMU [74] (AAAI’23) I3D 81.66 86.97
CLIP-TSA [17] (ICIP’23) ViT 82.19 87.58
VadCLIP [58] (AAAI’24) ViT 84.51 88.02
Yang et al. [62] (CVPR’24) ViT 83.68 87.79
Wu et al. [57] (CVPR’24) ViT 66.53 86.40
@@ 5.4 Qualitative Comparision
We provide qualitative comparisons between Holmes-VAU and existing MLLMs in Fig. 4. The results demonstrate that Holmes-VAU can accurately identify anomalies in videos and provide accurate and complete explanations, highlight the effectiveness and advantage of Holmes-VAU in perceiving video events and analyzing anomalies.
@@ 6 Conclusion
Acknowledgement This work is supported by the National Natural Science Foundation of China under grants U22B2053 and 623B2039, and in part by the Interdisciplinary Research Program of HUST (2024JCYJ034).
@@ A.2 Hierarchical Free-text Annotation
[linenos, frame=lines, breaklines]json "video": "v=2rfyeR-YaJw__#1_label_G-0-0", "n_frames": 1940, "fps": 24.0, "label": ["Explosion"], "clips": [[[5.583, 11.903], [11.903, 18.222], [18.222, 24.542]], [[36.167, 43.48], [43.48, 50.792]]], "clip_captions":[ [ "A military tank moving across a barren landscape with low-rise buildings and sparse vegetation. the sky is overcast, and the overall color palette is muted with earthy tones.", "A series of images depicting a barren landscape with a few buildings in the background. the foreground consists of a rocky terrain with sparse vegetation. the sky is overcast, and there are no visible people or moving objects.", "A silhouette of a person operating a large, mounted weapon on a rocky terrain under a clear sky. the individual appears to be adjusting or aiming the weapon." ], [ "A series of images depicting a hazy, dusty environment with buildings in the background. the focus is on a red light source that appears to be moving or flashing, possibly indicating some form of activity or event occurring in the area.", "A series of images depicting a military tank moving across a landscape with buildings in the background. the tank is seen from different angles, and there are visible explosions occurring in the distance." ] ], "events": [[5.583, 24.542], [36.167, 50.792]], "event_summary": [ "The anomaly exists, specifically an Explosion anomaly. The basis for judging the anomaly is the presence of an unexpected and sudden release of energy, likely accompanied by a loud noise and visible effects such as smoke, fire, or debris.", "The anomaly exists and its specific name is Explosion. The anomaly event is a series of visible explosions occurring in the distance as a military tank moves across a landscape with buildings in the background. The basis for judging the anomaly is the sudden and intense release of energy and light in the form of explosions, which is an unusual and notable event in the otherwise desolate and quiet war-torn environment depicted in the video"], "video_summary": "The anomaly exists, specifically named as Explosion. The anomaly event depicts a series of visible explosions occurring in the distance as a military tank moves across a landscape with buildings in the background, characterized by sudden and intense releases of energy and light, accompanied by loud noise and visible effects such as smoke, fire, or debris. The basis for judging the anomaly lies in the unusual and notable nature of these explosions, which stand out against the otherwise desolate and quiet war-torn environment depicted in the video, making them an unexpected and sudden release of energy that grabs attention."
@@ A.3 Hierarchical Instruction Data Construction
    "Describe the video briefly."
@@
    "Describe the main events that take place in this video."
@@
    "Give a short description of the video."
@@
    "What happened in this video?"
@@
    "Generate a brief caption for the video."
@@
    "Can you provide a brief description of the video?"
@@
    "Briefly describe the main subjects and their actions in the video."
@@
    "Provide a short overview of what happens in the video?"
@@
    "Describe the key moments that showcase the subjects’ activities in the video."
@@
    "Describe the sequence of events involving the main subjects in the video."
@@
    "What activities happen throughout the video?"
@@
    "Describe the main subjects and their roles in the video."
@@
    "What key moments stand out in the video?"
@@
    "What are the primary activities showcased in the video?"
@@
    "What happens to the main subjects as the video progresses?"
@@
    "What is a brief overview of what happens in the video?"
@@
    "Describe the main subjects and their contributions to the video."
@@
    "Describe the key events in the video."
@@
    "Describe the video’s main activities."
@@
    "Can you describe the main action in this video briefly?"
@@
    "Describe the video clip concisely."
@@
    "Provide a brief description of the given video clip."
@@
    "Summarize the visual content of the video clip."
@@
    "Give a short and clear explanation of the subsequent video clip."
@@
    "What types of anomalies are shown in the video clip?"
@@
    "Are there any anomaly events detected in the video?"
@@
    "Detect and classify the anomaly events in the video."
@@
    "Identify any abnormal behaviors depicted in the video."
@@
    "Determine whether there are anomaly events in the video and the specific name of the anomaly."
@@
    "What anomalies can be identified in the video?"
@@
    "What categories of anomalies can be found in the video?"
@@
    "Could you point out any abnormal actions in the video?"
@@
    "Point out the abnormal actions in the video."
@@
    "Are there anomalies observed in the video clip?"
@@
    "Describe the anomaly events observed in the video."
@@
    "Could you describe the anomaly events observed in the video?"
@@
    "Could you specify the anomaly events present in the video?"
@@
    "Give a description of the detected anomaly events in this video."
@@
    "Could you give a description of the anomaly events in the video?"
@@
    "Provide a summary of the anomaly events in the video."
@@
    "Could you provide a summary of the anomaly events in this video?""
@@
    "What details can you provide about the anomaly in the video?"
@@
    "How would you detail the anomaly events found in the video?"
@@
    "How would you describe the particular anomaly events in the video?"
@@
    "Why do you judge this event to be anomalous?"
@@
    "Can you provide the reasons for considering it anomalous?"
@@
    "Can you give the basis for your judgment of this event as an anomaly?"
@@
    "What led you to classify this event as an anomaly?"
@@
    "Could you provide the reasons for considering this event as abnormal?"
@@
    "What evidence do you have to support your judgment of this event as an anomaly?"
@@
    "Can you analyze the factors contributing to this anomalous event?"
@@
    "Could you share your analysis of the anomalous event?"
@@
    "What patterns did you observe that contributed to your conclusion about this event being an anomaly?"
@@
    "How do the characteristics of this event support its classification as an anomaly?"
@@
[linenos, frame=lines, breaklines]json "id": "Abuse007_x264_E0C2", "type": "clip", "video": "ucf-crime/clips/train/Abuse007_x264_E0C2.mp4", "conversations": [ "from": "human", "value": "Describe the key events in the video.", "from": "gpt", "value": "The man in shorts was slapped to the ground by the strong man opposite." ]
@@
[linenos, frame=lines, breaklines]json "id": "Abuse018_x264_E0", "type": "event", "video": "ucf-crime/events/train/Abuse018_x264_E0.mp4", "conversations": [ "from": "human", "value": "Identify any abnormal behaviors depicted in the video.", "from": "gpt", "value": "An anomaly exists, specifically an instance of Police Brutality.", "from": "human", "value": "Could you provide a summary of the anomaly events in this video?", "from": "gpt", "value": "The anomaly event involves the police officer kicking the man who rushed towards him, causing the man to fall heavily to the ground with his head on his back.", "from": "human", "value": "Could you share your analysis of the anomalous event?", "from": "gpt", "value": "The excessive use of force by the police officer, which deviates from standard protocol and may be considered an abuse of authority." ]
@@
[linenos, frame=lines, breaklines]json "id": "v=xLFbqPzSI2Y__#1_label_B4-0-0", "type": "video", "video": "xd-violence/videos/train/ v=xLFbqPzSI2Y__#1_label_B4-0-0.mp4", "conversations": [ "from": "human", "value": "Are there any anomaly events detected in the video?", "from": "gpt", "value": "The anomaly exists and its specific name is the Riot anomaly.", "from": "human", "value": "Could you give a description of the anomaly events in the video?", "from": "gpt", "value": "The anomaly event depicts a chaotic and intense urban unrest scene, starting with a chaotic street scene where individuals are running away from an area of fire and smoke, with people appearing to be in a state of panic or urgency, and some being sprayed with water by individuals in uniform. The scene escalates to show a growing fire in the background, police officers in riot gear standing on the street, and a formation of police officers spread out across the road, suggesting an ongoing riot or civil unrest. ", "from": "human", "value": "Can you give the basis for your judgment?", "from": "gpt", "value": "The presence of multiple indicators of a riot or emergency situation, including fire, smoke, and people fleeing in panic, as well as the sudden appearance of flames, smoke, and police in riot gear, and the presence of violent and destructive events, such as fires and explosions, and the response of authorities, including the use of tear gas or pepper spray, which deviate significantly from normal, peaceful urban activity." ]